OpenStudy (jenny.andrade):

Solve the system of equations. y = x^2 - 4 y = 2x - 1 Select one: a. (-2, 8) and (0, 4) b. (-1, -3) and (-3, 5) c. (-1, 3) and (3, -5) d. (-1, 3) and (0, 2)

OpenStudy (anonymous):

y = x^2 - 4, y = 2x - 1 So, you can write x^2 - 4 = 2x - 1 i.e x^2 -2x - 3 = 0 Now solve this quadratic equation.

OpenStudy (anonymous):

Can you solve it ?

OpenStudy (anonymous):

Factorization.

OpenStudy (anonymous):

What 2 numbers add up to -2 and multiply to -3?

OpenStudy (anonymous):

\[(X-? )(X+?) = 0\]

OpenStudy (cp9454):

correct option is not given

OpenStudy (zzr0ck3r):

or just plug in the answers and see what works:)

OpenStudy (anonymous):

Trial and error

OpenStudy (anonymous):

It is given.

OpenStudy (zzr0ck3r):

o god dont medal me for that, his was much better,

OpenStudy (cp9454):

then what is it @Johnbc

OpenStudy (anonymous):

Ha ha, it is great to know that you have a second method to solve this!

OpenStudy (anonymous):

@cp9454 Let the original poster learn to solve the problem.

OpenStudy (anonymous):

I'm sorry, but he's right. Plugging in doesn't seem to yield a correct answer.

OpenStudy (anonymous):

(x,y) and (x,y)

OpenStudy (anonymous):

I know, and none of them work for both sets.

OpenStudy (anonymous):

the closest one is answer b and the (-3,5) doesn't work for the second equation

OpenStudy (anonymous):

Most likely error in choices provided if they do not.

OpenStudy (anonymous):

So we can agree that there is no correct answer amongst the ones we are provided with. Right?